OpenStudy (kashif):

how under root of {2A^2 x (1+ cos theta)}=under root of {4A^2 x cos^2theta/2} ?

zepdrix (zepdrix):

Hey :) Welcome to OpenStudy! I'm not sure what your question is. There is no simpler way to express 1+cos(theta).

zepdrix (zepdrix):

\(\rm 1+\cos\theta\) is an expression, there is nothing to solve. Are you trying to solve this? \(\rm 1+\cos\theta=0\)

OpenStudy (freckles):

\[\cos(2 u)=\cos^2(u)-\sin^2(u) \\ \cos(2 u)=\cos^2(u)-(1-\cos^2(u)) \\ \cos(2 u)=2 \cos^2(u)-1 \\ \text{ solve for } \cos^2(u) \\ \cos^2(u)=\frac{1}{2}(\cos(2 u)+1) \\ \text{ Let } u=\frac{\theta}{2} \\ \cos^2(\frac{\theta}{2})=\frac{1}{2}(\cos( \theta)+1) \\ \text{ or let's write it as } \\ 2 \cos^2(\frac{\theta}{2})=\cos(\theta)+1 \\ \text{ so you can replace } \cos(\theta)+1 \text{ with } 2\cos^2(\frac{\theta}{2})\]

OpenStudy (freckles):

The identity is called the half angle identity.

OpenStudy (freckles):

\[\sqrt{2 A^2 (1+ \cos(\theta)} = \sqrt{2 A^2 (2 \cos^2(\frac{\theta}{2}))} =\sqrt{4 A^2\cos^2(\frac{\theta}{2})}\]

OpenStudy (freckles):

so both the left and right expressions are equal

OpenStudy (kashif):

thanks freckles

OpenStudy (freckles):

np

OpenStudy (freckles):

was this part of a calculus question you were doing?